OpenStudy (anonymous):

express sin 3theta in terms of sin theta

OpenStudy (anonymous):

sin\[\sin3\theta=3\sin \theta-4\sin ^{3}\theta\]

OpenStudy (anonymous):

could you please show me how you came about that? thanks